OpenStudy (cometailcane):

A sample of pure nitrogen gas is at 27 degrees C and 380 mm Hg occupied a volume of 492 mL. a. using the ideal gas law, determine the number of moles of gas in this sample? b. calculate the mass of the gas sample

OpenStudy (cuanchi):

a) \[P \times V = n \times R \times T\] \[n= \frac{ P \times V }{ R \times T }\] convert your pressure value from mmHg to atm with the conversion factor "1atm/760 mmHg) convert the temperature to K "K=C+273.15" convert the volume from mL to L (1L/1000mL) R= 0.082 atm L/mol K b) with the number of moles calculated from a) multiply by the MM of N2 (14 x 2) to get the mass of N2

OpenStudy (cometailcane):

thank you so much! the only question i have is on my paper it says R is gas constant (8.31 J/mol-K) and your saying its 0.082 atm L/mol K. which one should i use?

OpenStudy (cometailcane):

@Cuanchi

OpenStudy (surry99):

n must be in moles so what units must R be in? Write your units into your equation and you will see which units cancel out.

OpenStudy (cuanchi):

Yes! R= constant of the gases can have many different values depending on the units. It is like said the length of a book is 11 inches or said is 27.94 cm. For the general gas Law the most simple way to remember the R value is 0.082 atm L/mol K, because the other values in the formula are in the same units. Then for example if you want to express the R in place of atmospheres in mmHg, the value is going to be 0.082 x 760 = R= 62.32 mmHg L/mol K. In your case the 8.314 J is equivalent to 0.082 atm L/mol K. But the units in this case is Joule/mol K ; Joule = kg m^2/s^2= N x m = Pa x m^3. Atmospheres is unit of pressure, pressure is equal to force/area, force is mass times acceleration, acceleration is velocity over time. They are all derivative units. This form of R is mostly used in calculations of energy, since joules is a energy unit. You can check in Wikipedia they have a extensive list of values of R in different units http://en.wikipedia.org/wiki/Gas_constant

OpenStudy (cuanchi):

If you still want to use the R in J/mol K, you will have to convert the pressure to Pa=pascals (1 atmosphere =101 325 pascals) and the volume to m^3 (1m^3=1000L)